Mopeds (small motorcycles with an engine capacity below 50 cm3) are very popular in Europe because of their mobility, ease of operation, and low cost. Suppose the maximum speed of a moped is normally distributed with mean value 46.6 km/h and standard deviation 1.75 km/h. Consider randomly selecting a single such moped. (a) What is the probability that maximum speed is at most 50 km/h? (Round your answer to four decimal places.)

a. What is the probability that maximum speed is at most 50 km/h? (Round your answer to four decimal places.)

b. What is the probability that maximum speed is at least 47 km/h? (Round your answer to four decimal places.)

c. What is the probability that maximum speed differs from the mean value by at most 1.5 standard deviations? (Round your answer to four decimal places.)
Answer #1

ماتوقعه عدمعتها Xwn (46.6, 1:75). o pe X Ç so) = P(x-466 < 50-66-6) Tons 175 =plz c3 4 1.75 ) 20.9740 a pcx>47) = 1-Plu-46.6I PC Tu-11 21:50) pr-iis c2c1.S PC261.5) - Przc-1.5) = 0.86641 Ics Scanned with SI Cairnseamet